my moms age in years is the same as my brothers age in months.
My mothers age is 6 times my sisters age.
My age is the sum of my brother and sisters ages.
My fathers age in years is the number of seasons I've been alive.
My mother and father are the same age.

Answer by richwmiller(17219)

m=12b
m=6s
a=b+s
f=4a
m=f
Not enough info to pin it down to only one answer.
in terms of a
b = a/3, f = 4a, m = 4a, s = (2a)/3
in terms of b
a = 3b, f = 12b, m = 12b, s = 2b
If brother is 6 then
You are 18
father is 72
mother is 72
sister is 12
If brother is 2 then
You are 6
father is 24
mother is 24
sister is 4